Determine the net income of a “comparable” firm based on the following information: value of target firm = $8,000,000; net income of target firm = $400,000; stock price of “comparable” firm = $60.00; and 600,000 shares of stock outstanding for the comparable firm.
Answer #1

Value of comparable firm = price*shares = 60*600000=36000000

net income of comparable firm = net income of target firm*Value of comparable firm/value of target firm

=400000*36000000/8000000=1800000